Debbie's Cookies has a return on assets of 9.1 percent and a cost of equity of 12.2 percent. What is the pretax cost of debt if the debt–equity ratio is .87? Ignore taxes.
Answer #1

Debt-equity ratio=debt/equity

Hence debt=0.87 equity

Let equity be $x

Debt=$0.87x

Total=$1.87x

ROA=Respective costs*Respective weight

9.1=(x/1.87x*12.2)+(0.87x/1.87x*Cost of debt)

9.1=6.524064171+(0.87/1.87*Cost of debt)

Cost of debt=(9.1-6.524064171)*(1.87/0.87)

which is equal to

=5.54%(Approx).